Status Report

NASA Space Station On-Orbit Status 23 Jun 2004

By SpaceRef Editor
June 29, 2004
Filed under , , ,
NASA Space Station On-Orbit Status 23 Jun 2004

All ISS systems continue to function nominally except those noted previously or below. 

After their very successful Orlan dry-run activities yesterday, the crew is now in the home stretch for tomorrow’s EVA evening, scheduled to begin at 5:50pm EDT with DC1 EV1 hatch opening.

On the current sleep/wake cycle, the crew’s wakeup time today was shifted by six hours to 8:00am EDT, with bedtime tonight at 10:30pm.  Tomorrow, wakeup is set for 9:10am and sleep for 5:00am the next day. By the weekend, the schedule will have returned to the usual pattern.

CDR Padalka and FE/SO Fincke restored their Orlan-M suits to readiness for the spacewalk by filling the DIDBs (disposable in-suit drink bags) and installing them in the suits.  Padalka also retrieved a Fresnel lens from stowage and mounted it in his Orlan, which had lacked the viewing aid.

Gennady equipped both Orlans with “Pille-MKS” radiation sensors, which he removed from their exposure locations in the RS after recording their dose measurements.  A third sensor, for background readings, was placed in the SM cabin.   [“Pille” has ten sensors normally situated at various locations in the RS (port cabin window, stbd cabin window, ASU toilet facility, control panel, etc.).  Dosage values are called down or downlinked via Regul Paket/Email or OCA.]

The crew had another review of the final spacewalk timeline and later tonight will tag up with EVA specialists at TsUP/Moscow to discuss the venture via S-band S/G (space-to-ground).   [After egressing the DC1 at ~5:56pm EDT, the first 30-35 minutes will be spent on preparing the GStM-1 “Strela” cargo boom, mounted on the DC1 hull, and extending it (by manually cranking its pitch and yaw drive handles) towards the FGB, where its end effector (TU) will be attached at a handrail (#1076).  The crane, measuring 3 m retracted and 15 m in fully extended configuration, then serves as handrail and attachment for the crew’s safety tethers until U.S. tethers are secured at the PMA and later on the S0 truss.  Removal and replacement (R&R) of the failed RPCM, to start at or about 8:30pm, requires prior opening of the door of the portside power assembly SPDA (secondary power distribution assembly).  After the R&R and the restoration of external configurations, including retraction and securing of the GStM-1, ingress of the last spacewalker, Padalka (EV1), is expected at ~11:30pm.  The two Mission Control Centers will take turn in leading the spacewalk, with TsUP in charge of Russian segment (RS) operations and egress, translation, return, ingress and any Russian task activities, while MCC-Houston has the lead for detailed ops dealing with the U.S. segment, the RPCM R&R and the power-down of the external Lab DDCU (dc-to-dc converter unit) 2A power source of the RPCM prior to the R&R, and its subsequent power-up.]

The CDR deactivated the IK0501 gas analyzer (GA) in the SM and exchanged its BF carbon dioxide (CO2) filter assembly with a new unit from the FGB (replaced last: 5/14).  GA was reactivated and the spent BF stowed for return to Earth.  [IK0501 is an automated system for measuring CO2, O2, and H2O in the air as well as the flow rate of the gas being analyzed.]

In preparation for the EVA, which will require the station reconfigured to “unmanned mode”, Gennady powered down the new “Sputnik-SM” ham radio system in the SM.

Later tonight, Padalka will install an additional portable air repress bottle (BNP) #9 in the DC1 docking module to prepare it for the EVA, ensuring adequate availability of air to repress the airlock.

The FE conducted the daily routine maintenance on the SOZh life support system, today including the weekly BRPK air/condensate water separator system inspection, and conducted the regular status checkup of the autonomous PCG-STES010 payload in the Lab.

To prepare the Ku-band TV system for covering the EVA, Mike connected the UOP (utility outlet panel) bypass power cable to the RWS DCP (robotics workstation display & control panel).  He then configured the Sony PD100 camcorders in the Node and Lab for situational (internal) monitoring during the EVA.

Since hatches will be closed between the Russian and U.S. segments (and also between the SM and the DC1), later tonight Fincke will have to reconfigure the OpsLAN onboard laptop network appropriately.  He will also attempt to set up a ThinkPad A31p NGL (Next Generation Laptop) for unattended monitoring of the EVA from the Lab via its NetMeeting application.   [The experimental setup will be tested after the OpsLAN reconfiguration for hatch closure has been completed.]

The FE took the weekly readings for monitoring carbon monoxide (CO) and hydrogen chloride (HCl) contamination remaining in the new CSA-CPs (compound specific analyzer-combustion products) #1015 and #1016, which are still in their normal decontamination period (their oxygen sensors can be used for O2 monitoring, however).   [The older CSA-CP #1010, with nonfunctioning O2 sensor, is used for combustion products monitoring, and CSA-CP #1003 was stowed because its O2 sensor has exceeded its calibration period.]

Later tonight, Gennady is scheduled to terminate the current run of the Molniya-SM/LSO (GFI-10) geophysical experiment started on 6/20, by first deactivating the two French computers EGE 1 and EGE 2, the LSO 06 instrument and the Russian RBS power outlet, then disassembling the equipment and stowing it.   [Molniya-SM/LSO automatically records storm phenomena and other related events in the Earth’s equatorial regions.  The experiment is controlled from the EGE-1 laptop, loaded with orbital sighting predictions using an up-to-date NORAD tracking TLE (two-line element) provided by NASA.  Objective of LSO was to study rare optical phenomena occurring in the upper layers of Earth’s atmosphere, so-called “sprites” (i.e., puzzling glow phenomena observed above thunderstorm clouds).  LSO was originally part of Claudie Haigneré’s French “Andromeda” payload package of taxi mission 3S that could not be performed as planned during Increment 4 due to an ISS flight attitude conflict.]

Shortly before sleep time tonight, the CDR will break out the “Urolux” equipment, setting it up for the Russian biochemical urine test (PZE MO-9), a standard requirement before and after Orlan-suited activities.   [The MO-9 analysis uses the sophisticated in-vitro diagnostic apparatus “Urolux” developed originally for the Mir program.]

Both crewmembers are scheduled to work out on TVIS, RED exerciser, and VELO cycle with load trainer in their daily 2.5-hr. physical exercise program.

Ground-commanded preparations for the EVA included truss power reconfigurations for the RCPM R&R, such as swapping outside pump module heater strings for the NTA (nitrogen tank assembly) and terminating a power-on test of the loop B PCVP (pump & control valve package) because of tomorrow’s LA2A DDCU deactivation before the R&R.
Today’s CEO (Crew Earth Observations) photo targets, in the current LVLH attitude no longer limited by flight rule constraints on the use of the Lab nadir/science window, except for the shutter closure and condensation-prevention plan (limited to 90 min. in 24 hours), were Navassa Island reefs, Caribbean (looking a touch left.  ISS passed close to the east end of Jamaica which was clear), and Andes dust (Dynamic event.  Winter is the time for dust plumes to be raised by westerly winds on the high desert of the central Andes.  As the present front approached and winds picked up, the crew was to shoot any dust that could be seen.  Plumes can be small, restricted to the high desert; or they can be large and extend hundreds of miles eastward over the Argentine low country.  These dust plumes have high geological interest but are hard to see on weather satellite images.  The crew may have seen nothing, but conditions were right.

CEO images can be viewed at the websites.

See also the website “Space Station Challenge” at

U.S. & Russian Segment Status  (as of today, 1:50pm EDT)

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is On.  Vozdukh CO2 scrubber is On.  U.S. CDRA CO2 scrubber is Off.  TCCS (trace contaminant control subsystem) is operating.  SM Gas Analyzer has been calibrated and is used for ppO2 & ppCO2 monitoring.  MCA (major constituents analyzer) is in Life Extending Mode (LEM).  BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode.  RS air conditioner SKV-1 is On, SKV-2 is Off (SM panel mods completed 4/8; SKV-2 activation failed 4/20; currently inoperable due to leaky condensate hose).  SFOG slot #2 fan suspect (not usable).
  • SM Working Compartment:  Pressure (mmHg) — 742; temperature (deg C) — 25.5; ppO2 (mmHg) — 154.2; ppCO2 (mmHg) — 3.7.
  • SM Transfer Compartment:  Pressure (mmHg) — 751; temperature (deg C) — 20.0.
  • FGB Cabin:  Pressure (mmHg) — 756; temperature (deg C) — 22.0.
  • Node:  Pressure (mmHg) — 750.25; temperature (deg C) — 23.5 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • U.S. Lab:  Pressure (mmHg) — 752.19; temperature (deg C) — 24.8; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • Joint Airlock (Equip. Lock):   Pressure (mmHg) — 752.39; temperature (deg C) — 22.6; shell heater temp (deg C) — n/a, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • (n/a = data not available)
  • PMA-1:  Shell heater temp (deg C) — n/a
  • PMA-2:  Shell heater temp (deg C) — n/a.

Electrical Power Systems (EPS):

  • Both P6 channels fully operational.  BGA (beta gimbal assembly) 2B and 4B both in directed position (Blind mode, non solar-tracking).
  • SM batteries:  All batteries (8) are on line in “Partial Charge” mode.  
  • FGB batteries:  Battery #5 is off line; all other batteries (5) are on line in “Partial Charge” mode.
  • Plasma Contactor Unit PCU-1 is in Standby mode; PCU-2 is in Standby mode.

Command & Data Handling Systems:

  • C&C-1 MDM is prime, C&C-2 is backup, and C&C-3 is in standby.
  • GNC-2 MDM is prime; GNC-1 is backup.
  • INT-2 is operating; INT-1 is Off.
  • EXT-1 is On (primary), EXT-2 is Off (backup).
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-1 MDM is Off; PL-2 MDM is Operational.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational (string 1 dropped out 11/22/03).
  • SM Central Computer (TsVM): 2 redundant lanes (of 3) operational (string #3 dropped out 10/22/03).
  • FGB MDM-1 is Off (failed, 11/21/03); MDM-2 is Operational.

Propulsion System:

  • Total propellant load available: 4002 kg (8823 lb) as of 6/10/04;  [SM(552) + FGB(2811) + Progress M(639)].  (Capabilities: SM — 860 kg; FGB — 6120 kg).

Attitude Control Systems:

  • 2 CMGs on-line (CMG-1 failed, since 6/6/02; CMG-2’s RPC-17 failed 4/21/04).
  • State vector source — U.S. SIGI-1 (GPS)
  • Attitude source — U.S. SIGI-1 (GPS)
  • Angular rate source — RGA-1

Flight Attitude:

  • LVLH XVV (local vertical/local horizontal = “earth-fixed”: z-axis in local vertical, x-axis in velocity vector [yaw: -10 deg, pitch: -9 deg, roll: 0 deg]), with CMG/TA (thruster assist) Momentum Management, until 6/28, following the EVA.

Communications & Tracking Systems:

  • FGB MDM-1 is powered Off; FGB MDM-2 is operational.
  • All other Russian communications & tracking systems are nominal.
  • S-band is operating nominally (on string 2).
  • Ku-band is operating nominally (may require a mask).
  • Audio subsystem is operating nominally (IAC-1 is prime, IAC-2 is off).
  • Video subsystem operating nominally.
  • HCOR (high-rate communications outage recorder) is operating nominally.


  • SSRMS/Canadarm2 based at Lab PDGF/LEE A, operational on redundant string, off on prime.
  • MBS: KA (keep alive) power on both strings. 
  • MT: latched and mated at WS4. 
  • POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.

ISS Orbit  (as of this morning, 7:09am EDT [= epoch]):

  • Mean altitude — 361.0 km
  • Apogee — 364.6 km
  • Perigee — 357.4 km
  • Period — 91.8 min.
  • Inclination (to Equator) —  51.6333 deg
  • Eccentricity — 0.0005329
  • Solar Beta Angle — -1.7 deg
  • Orbits per 24-hr. day — 15.69
  • Mean altitude loss in last 24 hours — 80 m
  • Revolutions since FGB/Zarya launch (Nov. ’98)  — 31930

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see

SpaceRef staff editor.